OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (7, 4); Slope: 6 y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7)

OpenStudy (anonymous):

y-4=6(x-7)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

You put the x with the x and the y with the y. And so the general equation is y - __ = m(x - __)

OpenStudy (anonymous):

thx